# STGFormer: Spatio-Temporal GraphFormer for 3D Human Pose Estimation in Video

### v1

@@ III-B Overall Architecture
Joint-based embedding. Given a 2D pose sequence represented as $\mathbf{P}_{2D}\in\mathbb{R}^{T\times N\times 2}$, with $T$ and $N$ signifying the number of frames and body joints per frame respectively, we initially map $\mathbf{P}_{2D}$ to high-dimensional embeddings using a joint-based embedding layer. This layer independently processes each 2D coordinate by applying a fully connected layer, followed by a GELU activation. Consequently, the joint-based embedding layer generates features of dimension $T\times N\times F$. It is important to note that in the previous transformer, the embedding layer consolidates all joint coordinates in each frame into a singular vector. This approach reduces the computational demands of subsequent transformer blocks, yet it compromises spatial discrimination. In contrast, our method retains the spatial dimension $N$, while still reducing computational costs through the use of spatio-temporal criss-cross attention.
Spatio-Temporal GraphFormer. The STGFormer block is derived from the traditional transformer block, replacing the original multi-head self-attention layer with spatio-temporal criss-cross graph attention (STG attention). Additionally, a new dual-path modulated hop-wise regular GCN (MHR-GCN) is integrated into the STGFormer block to enhance the descriptive capability of local structures. Detailed discussions on the STG attention and MHR-GCN are provided in Sections III-C and III-D, respectively.
Regression head. A linear regression head is implemented after the STG blocks to predict the 3D pose coordinates $\hat{\mathbf{P}}_{3D}\in\mathbb{R}^{T\times N\times 3}$. The entire architecture is optimized by minimizing the Mean Squared Error (MSE) between the predicted 3D coordinates $\mathbf{\hat{P}}_{3D}$ and the ground-truth 3D coordinates $\mathbf{P}_{3D}$:
@@
III-C Spatio-temporal Criss-cross Graph Attention
Our attention mechanism is specifically designed to effectively model the dependencies between joints. Inspired by [25], we capture spatial and temporal contexts in parallel across different channels, thereby avoiding the quadratic computational cost associated with full attention. Specifically, our mechanism encodes spatio-temporal graph structural relationships into the attention layer to capture structural information between nodes. While modeling the spatial and temporal relationships of the human body, a learnable embedding is allocated based on their structural relationships.
To provide clarity, as shown in figure 1(b), the input embedding $\mathbf{X}\in\mathbb{R}^{T\times N\times F}$ is initially mapped to queries $\mathbf{Q}\in\mathbb{R}^{T\times N\times F}$, keys $\mathbf{K}\in\mathbb{R}^{T\times N\times F}$, and values $\mathbf{V}\in\mathbb{R}^{T\times N\times F}$, all sharing the same dimensions. These are subsequently split into two equal groups along the channel dimension. For clarity in notation, we refer to these split matrices as the time group $\{\mathbf{Q}_{T},\mathbf{K}_{T},\mathbf{V}_{T}\}$ and the space group $\{\mathbf{Q}_{S},\mathbf{K}_{S},\mathbf{V}_{S}\}$. Following this, separate self-attention modules are utilized to calculate the temporal and spatial correlations.
Previous models necessitated the explicit specification of distinct positions or the encoding of positional dependencies within layers. For sequential data, embeddings can be assigned to each position (i.e., absolute position encoding [37]) or encode the relative distance between any two positions in Transformer layers (i.e., relative position encoding [38]). However, for graphs, nodes are not sequentially arranged but are located in multidimensional spaces and connected by edges. To effectively incorporate the structural information of graphs into the model, we introduce a novel spatial encoding technique.
Specifically, we consider a function that measures the spatio-temporal relationships among elements within the graph, which is defined through the connectivity between nodes. We depict the temporal and spatial connectivity relationships of the human body structure in continuous sequences through graphs, and encode these relationships as inputs for spatio-temporal attention on multiple channels in parallel. For temporal encoding and spatial encoding, they are each respectively represented as follows:
@@ III-C Spatio-temporal Criss-cross Graph Attention
where ${b^{t}}_{\psi(v_{i},v_{j})}$ and ${b^{s}}_{\phi(u_{i},u_{j})}$ are learnable scalars, each indexed by their respective functions $\psi(v_{i},v_{j})$ and $\phi(u_{i},u_{j})$.
Consequently, the above two correlation modules operate in parallel and adhere to the self-attention scheme for feature contextualization. They calculate token-to-token affinities by contextualizing from specific axial perspectives, and effectively complement each other. Finally, we concatenate the outputs from both attention layers along the channel dimension:
@@
where $concat$ performs the concatenation. The resultant receptive field of STGA resembles a criss-cross of spatial and temporal axes, and stacking multiple STGA blocks can approximate complete spatio-temporal attention.
III-D Dual-path Modulated Hop-wise Regular GCN Module
@@
To advance the understanding of spatio-temporal relationships among data by directly learning their interconnections, this paper introduces a novel GCN module. Our dual-path modulated hop-wise regular GCN module splits the spatio-temporal graph convolution to independently learn their features, capturing high-order connections and long-range dependencies through an advanced hop-wise approach. Moreover, it employs residual connections to mitigate gradient vanishing and facilitates iterative reuse of features.
As shown in Figure 2, we separate the temporal and spatial graph learning into parallel paths to minimize interference between time and space dimensions. In spatial modulation, a replica of the graph is created for each time step, while in temporal modulation, edges representing temporal relations are introduced to capture dependencies across time, effectively integrating spatial and temporal information.
$\tilde{\mathbf{H}}_{j}^{(\ell)}=$ $\tilde{\mathbf{A}}_{j}(({\mathbf{H}_{S}}^{(\ell)}\mathbf{W}_{j}^{(\ell)})\odot\mathbf{M}_{j}^{(\ell)})$ (13)
$\tilde{\mathbf{H}}_{k}^{(\ell)}=$ $\tilde{\mathbf{A}}_{k}(({\mathbf{H}_{T}}^{(\ell)}\mathbf{W}_{k}^{(\ell)})\odot\mathbf{M}_{k}^{(\ell)})$
@@
where $\mathbf{A}_{j}$ and $\mathbf{A}_{k}$ are the spatial and temporal adjacency matrices, respectively, and $\mathbf{W}_{j}^{(\ell)}$, $\mathbf{W}_{k}^{(\ell)}$ are the corresponding learnable weight matrices. The adjacency modulation matrices $\mathbf{M}_{k}^{(\ell)}$ and $\mathbf{M}_{j}^{(\ell)}$ allow the model presented in this paper to respond flexibly to data variations.
@@
As illustrated in Figure 3, the concept of hop-wise is introduced in MHR-GCN to capture high-order connection information and long-range dependencies. We use k-hop temporal and j-hop spatial neighbors as shown in figure 1(c). Inspired by the RSNet [13], residual connections are employed in both the temporal and spatial paths to prevent gradient vanishing and bolster feature reuse. The spatio-temporal layer-wise propagation rule of the regular dual-path hop-wise graph network with weight and adjacency modulation is given by
@@
where $\mathbf{X}_{S}{\tilde{\mathbf{W}}_{j}}^{(\ell)}$ and $\mathbf{X}_{T}{\tilde{\mathbf{W}}_{k}}^{(\ell)}$ are the regular connection, $\sigma\left(\cdot\right)$ is an element-wise nonlinear activation function such as the Gaussian Error Linear Unit (GELU), $\boxplus$ denotes concatenation, ${\tilde{\mathbf{W}}_{j}}^{(\ell)}\in\mathbb{R}^{F\times F}$ and ${\tilde{\mathbf{W}}_{k}}^{(\ell)}\in\mathbb{R}^{F\times F}$ are the learnable weight matrices, $\mathbf{H}_{S}^{(\ell)}\in\mathbb{R}^{N\times F}$ and $\mathbf{H}_{T}^{(\ell)}\in\mathbb{R}^{N\times F}$ are the input feature matrix of the $\ell$-th layer, $\mathbf{H}_{S}^{(\ell+1)}\in\mathbb{R}^{N\times F}$, $\mathbf{H}_{T}^{(\ell+1)}\in\mathbb{R}^{N\times F}$ are the output feature matrix. Thus, the equation of the $\ell+1$ layer feature is reformulated as
@@
IV Experiments
In this section, we first introduce the two benchmark datasets i.e., Human3.6M [39] and MPI-INF-3DHP [40], along with the evaluation metrics used to assess the proposed framework and the experimental setup. Subsequently, we evaluate the proposed framework and compare it with the current state-of-the-art methods, presenting both quantitative and qualitative results. Finally, we verify the impact of each proposed design through ablation studies.
IV-A Experimental Setup
@@
Human3.6M [39] is the most widely used dataset in the evaluation of 3D human pose estimation. It encompasses a vast collection of 3.6 million poses captured using RGB and ToF cameras from diverse viewpoints within a real-world setting. Adhering to the widely used standard protocol, we set five subjects (S1, S5, S6, S7, and S8) for training, and two subjects (S9 and S11) for evaluation. We use Mean Per Joint Position Error (MPJPE) to measure the error under two protocols: Protocol 1 (referred to as P1) and Protocol 2 (referred to as P2). P1 involves calculating MPJPE between the estimated pose and the ground truth after aligning their root joints; P2 involves computing Procrustes-MPJPE (PA-MPJPE), where the ground truth and the pose prediction are further aligned through a rigid transformation.
MPI-INF-3DHP [40] is a large-scale dataset that offers over 2K videos featuring joint annotations of 13 keypoints in outdoor scenes, suitable for 2D and 3D human pose estimation. The ground truth was obtained using a multi-camera arrangement and a marker-less MoCap system. Here, we report the MPJPE, Percentage of Correct Keypoints (PCK) at a threshold of 150mm, and Area Under the Curve (AUC) results as the evaluation metrics.
@@
We evaluate the performance of our model by comparing it with state-of-the-art video-based 3D human pose estimation methods, including transformer-based approaches such as StridedFormer [22], MHFormer [21], MixSTE [20], P-STMO [41], HSTFormer [18], Einfalt et al. [19], and PoseFormerV2 [16], as well as GCN-based methods like UGCN [42] and GLA-GCN [11]. Additionally, we also compare our model with methods that combine graphs and transformers, such as STCFormer [25] and MotionAGFormer [43].
@@
In the experiments, we utilize two kinds of 2D pose sequences as input including the estimated 2D pose by the pre-trained CPN [44] and the real 2D pose (ground truth). For model training, the network parameters are optimized for 40 epochs using the Adam optimizer, with a batch size of 128. The basic learning rate is 0.001, with a decay of 0.97 per epoch. We instantiate our architecture with L = 6 for transformer layers, F = 256 for feature embedding dimensions, H = 8 for self-attention heads, N = 2 for MHR-GCN layers. Our model is implemented using the PyTorch toolkit, version 2.1.0, and runs on RTX 4090 GPU with driver version 525.105.17 and CUDA version 12.0.
IV-B Results and Analysis

### v2

@@ 3.2 Overall Architecture
Joint-based embedding.
Given a 2D pose sequence represented as $\mathbf{P}_{2D}\in\mathbb{R}^{T\times N\times 2}$, where $T$ denotes the number of frames and $N$ denotes the number of body joints per frame, we first map $\mathbf{P}_{2D}$ to high-dimensional embeddings using a joint-based embedding layer. This layer independently processes each 2D coordinate by applying a fully connected layer, followed by a GELU activation. As a result, the joint-based embedding layer produces features with dimensions $T\times N\times F$.
Spatio-Temporal GraphFormer.
The STGFormer block is derived from the transformer block in Eq.(1), where the original multi-head self-attention layer is replaced with spatiotemporal criss-cross graph attention. Additionally, a novel dual-path MHR-GCN is incorporated into the STGFormer block to enhance its ability to model local structures. Detailed discussions on the STG attention and dual-path MHR-GCN are provided in Sections 3.3 and 3.4, respectively.
Regression head.
A linear regression head follows the STGFormer blocks to predict the 3D pose coordinates $\hat{\mathbf{P}}_{3D}\in\mathbb{R}^{T\times N\times 3}$. The entire architecture is optimized by minimizing the Mean Squared Error (MSE) between the predicted 3D coordinates $\hat{\mathbf{P}}_{3D}$ and the ground-truth 3D coordinates $\mathbf{P}_{3D}$:
@@
3.3 Spatio-temporal Criss-cross Graph Attention
Our attention mechanism is specifically designed to leverage prior knowledge of the human skeleton more effectively, ultimately improving performance in 3D human pose estimation. Inspired by [25], we capture spatial and temporal contexts in parallel across different channels, thereby avoiding the quadratic computational cost associated with full attention. Specifically, our mechanism encodes spatio-temporal graph structural relationships into the attention layer to capture structural information between nodes. While modeling the spatial and temporal relationships of the human body, a learnable embedding is assigned based on their structural relationships.
For clarity, as shown in Figure 1(b), the input embedding $\mathbf{X}\in\mathbb{R}^{T\times N\times F}$ is first mapped to queries $\mathbf{Q}\in\mathbb{R}^{T\times N\times F}$, keys $\mathbf{K}\in\mathbb{R}^{T\times N\times F}$, and values $\mathbf{V}\in\mathbb{R}^{T\times N\times F}$, all sharing the same dimensions. These are subsequently split into two equal groups along the channel dimension. We refer to these split matrices as the time group $\{\mathbf{Q}_{T},\mathbf{K}_{T},\mathbf{V}_{T}\}$ and the space group $\{\mathbf{Q}_{S},\mathbf{K}_{S},\mathbf{V}_{S}\}$. Following this, separate self-attention modules are utilized to calculate the temporal and spatial correlations. Previous models required the explicit specification of distinct positions or the encoding of positional dependencies within layers. For sequential data, embeddings can be assigned to each position (i.e., absolute position encoding [38]) or encode the relative distance between any two positions in Transformer layers (i.e., relative position encoding [39]). However, for graphs, nodes are not sequentially arranged but are positioned in multidimensional spaces and connected by edges.
To further integrate edge features into the attention layer and incorporate the structural information of graphs into the model, we propose a novel encoding technique in STGFormer. The temporal and spatial relationships of human joints are encoded in the softmax attention, which allows the model to accurately capture spatial dependencies in the graph. Specifically, for each pair of nodes, we compute their connectivity with other nodes and then incorporate this information into the attention module. Equipped with these encodings, STGFormer is able to better model pairwise node relationships and represent the graph. This can be viewed as a refinement of the basic attention scores, endowing the model with the ability to distinguish different relative spatiotemporal positions and thereby better capture prior knowledge associated with spatial distance and temporal adjacency. For temporal encoding and spatial encoding, they are each respectively represented as follows:
@@ 3.3 Spatio-temporal Criss-cross Graph Attention
where $A_{ij}^{t}$ and $A_{ij}^{s}$ represent the attention scores between the $i$-th and $j$-th nodes in the temporal and spatial attention modules, respectively; $h_{i}^{t}$, $h_{j}^{t}$, $h_{i}^{s}$ and $h_{j}^{s}$ are the feature representation vectors of the $i$-th and $j$-th nodes corresponding to the temporal and spatial dimensions; $W_{Q}^{t}$, $W_{K}^{t}$, $W_{Q}^{s}$ and $W_{K}^{s}$ denote the learnable projection matrices that map the representations into the query and key spaces in the temporal and spatial attention branches; $\sqrt{d}$ represents the scaling factor; and ${b^{t}}_{\psi(v_{i},v_{j})}$ and ${b^{s}}_{\phi(u_{i},u_{j})}$ are learnable scalars, each indexed by its respective function $\psi(v_{i},v_{j})$ or $\phi(u_{i},u_{j})$.
Here we discuss the advantages of our proposed method. We introduce a function $b(\cdot)$ that quantifies the spatio-temporal relationships among elements within the graph, which is defined through the connectivity between nodes. Specifically, we depict the temporal and spatial connectivity relationships of the human body structure in continuous sequences through graphs, and encode these relationships as inputs for spatio-temporal attention on multiple channels in parallel. For example, if ${b}_{\psi(v_{i},v_{j})}$ is learned to be a decreasing function of $\psi(v_{i},v_{j})$, the model is more inclined to assign higher attention to nodes in close proximity and lower attention to those that are farther away.
Consequently, the two correlation modules described above operate in parallel and adhere to the self-attention scheme for feature contextualization. They calculate token-to-token affinities by contextualizing from specific axial perspectives, and effectively complement each other. Finally, we concatenate the outputs from both attention layers along the channel dimension:
@@
where $concat$ performs the concatenation. The resultant receptive field of our attention resembles a criss-cross of spatial and temporal axes, and stacking multiple STG attention blocks approximates complete spatio-temporal attention.
3.4 Dual-path Modulated Hop-wise Regular GCN Module
@@
To improve our understanding of spatiotemporal graph relationships by directly learning their interconnections, this paper introduces a novel GCN module. Our dual-path modulated hop-wise regular GCN module splits spatiotemporal graph convolution to independently learn features, thereby reducing computational complexity and preserving both temporally dynamic features and original (or high-dimensional) representations of spatial structures. The parallel dual-path strategy avoids the sequential processing typical of serial architectures, thereby mitigating the risk of prematurely compressing certain features. In addition, it employs a spatiotemporal hop-wise approach to capture higher-order dependencies and incorporates a residual connection to mitigate gradient vanishing and facilitate the iterative reuse of features.
Consider a graph convolutional layer in the vanilla GCN [40] that transforms and aggregates input features according to the following equation:
$\mathbf{H}^{\prime}=\sigma\left(\mathbf{WH}\tilde{\mathbf{A}}\right)$ (6)
where $\mathbf{H}^{\prime}$ is the updated feature matrix, $\sigma(\cdot)$ represents the activation function, $\mathbf{W}$ is the weight matrix, $\mathbf{H}$ is the input feature matrix, and $\tilde{\mathbf{A}}$ is the adjacency matrix. The $i$-th row of the output feature matrix can be expressed as:
$\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}{\mathbf{Wh}_{j}}\tilde{a}_{ij}\right)$ (7)
where $\mathbf{h}_{i}^{\prime}$ is the updated feature vector of node $i$, $\mathbf{h}_{j}$ is the input feature vector of node $j$, $\mathcal{N}_{i}$ denotes the set of neighbors of node $j$, $\tilde{a}_{ij}$ is the element of the modified adjacency matrix corresponding to nodes $i$ and $j$. However, the shared feature transformation weight in the vanilla GCN limits its ability to learn diverse relational patterns from different body joint nodes. Liu et al. [26] address this limitation by utilizing a unique weight matrix $\mathbf{W}_{j}\in\mathbb{R}^{F^{\prime}\times F}$ for each node $j$. Nevertheless, as the number of joints and frames grows, this approach leads to a significant rise in the total number of parameters. To reduce model size, a weight modulation strategy [27] is introduced, which applies a shared transformation but modulates it differently for each node, as expressed by the following equation:
$\mathbf{h}_{i}^{\prime}=\sigma\left(\sum_{j\in\mathcal{N}_{i}}{(}\mathbf{m}_{j}\odot\mathbf{W})\mathbf{h}_{j}\tilde{a}_{ij}\right)$ (8)
where $\mathbf{m}_{j}$ is a learnable modulation vector for each neighboring node $j$ and $\odot$ denotes element-wise multiplication.
As illustrated in Figure 2, we separate the temporal and spatial graph learning components into parallel pathways in our MHR-GCN. This decoupling mitigates the risk of feature entanglement and allows for independent optimization and regularization, leading to a more efficient and robust representation. Furthermore, by subsequently merging these two pathways, the model leverages complementary information that collectively enhances performance in tasks demanding detailed spatiotemporal understanding. In spatial modulation, a replica of the graph is created for each time step, while in temporal modulation, edges representing temporal relations are introduced to capture dependencies across time. As a result, by focusing on local spatial structures within individual frames and replicating the graph at every time step, the model can track variations in node connectivity and other spatial features with greater accuracy. The updated feature representations at layer $\ell$ for the spatial node $j$ and the temporal node $k$ can be expressed as follows:
$\tilde{\mathbf{H}}_{j}^{(\ell)}=$ $\sigma\left(\tilde{\mathbf{A}}_{j}(({\mathbf{H}_{S}^{(\ell)}}\mathbf{W}_{j}^{(\ell)})\odot\mathbf{M}_{j}^{(\ell)})\right)$ (9)
$\tilde{\mathbf{H}}_{k}^{(\ell)}=$ $\sigma\left(\tilde{\mathbf{A}}_{k}(({\mathbf{H}_{T}^{(\ell)}}\mathbf{W}_{k}^{(\ell)})\odot\mathbf{M}_{k}^{(\ell)})\right)$
@@
where $\tilde{\mathbf{A}}_{j}$ and $\tilde{\mathbf{A}}_{k}$ are the spatial and temporal adjacency matrices, ${\mathbf{H}_{S}^{(\ell)}}$ and ${\mathbf{H}_{T}^{(\ell)}}$ are the input feature representations at layer $\ell$ for the spatial and temporal streams, respectively, and $\mathbf{W}_{j}^{(\ell)}$, $\mathbf{W}_{k}^{(\ell)}$ are the corresponding learnable weight matrices. The adjacency modulation matrices $\mathbf{M}_{k}^{(\ell)}$ and $\mathbf{M}_{j}^{(\ell)}$ allow the model presented in this paper to respond flexibly to data variations.
@@
With increasing network depth, the model becomes more susceptible to suboptimal convergence (i.e., falling into local optima), and the resulting node representations tend to become overly smooth, making it difficult to distinguish features across different nodes. To address this limitation, we employ residual regular connections [13] across layers to facilitate feature reuse in both the temporal and spatial paths. Moreover, as illustrated in Figure 3, the concept of hop-wise is introduced in MHR-GCN to capture high-order connection information and long-range dependencies. In this paper, we utilize $k$-hop temporal and $j$-hop spatial neighbors, as depicted in Figure 1(c), to further expand the receptive field.
Finally, the spatio-temporal layer-wise propagation rule of the regular dual-path hop-wise graph network with weight and adjacency modulation is formulated as follows, where $\mathbf{H}_{S}^{(\ell)}\in\mathbb{R}^{N\times F}$ and $\mathbf{H}_{T}^{(\ell)}\in\mathbb{R}^{N\times F}$ represent as the input feature matrices at the $\ell$-th layer:
@@
where $J$ and $K$ represent the hop range considered in the spatial and temporal domain, $\tilde{\mathbf{H}}_{j}^{(\ell)}$ and $\tilde{\mathbf{H}}_{k}^{(\ell)}$ are the intermediate representations obtained at layer $\ell$ through $j$-hop or $k$-hop neighbor aggregation, $\mathbf{X}_{S}{\tilde{\mathbf{W}}_{j}}^{(\ell)}$ and $\mathbf{X}_{T}{\tilde{\mathbf{W}}_{k}}^{(\ell)}$ are the regular connection, $\mathbf{H}_{S}^{(\ell+1)}\in\mathbb{R}^{N\times F}$, $\mathbf{H}_{T}^{(\ell+1)}\in\mathbb{R}^{N\times F}$ are the output feature matrix, $\boxplus$ denotes concatenation. The output feature vectors $\mathbf{h}_{p}^{(\ell+1)}$ and $\mathbf{h}_{u}^{(\ell+1)}$ of node $p$, $u$ on the spatial and temporal paths can be represented as follows:
$\mathbf{h}_{p}^{(\ell+1)}=$ $\sigma\left(\sum_{q\in\mathcal{N}_{p}}{\tilde{a}_{pq}}\mathbf{h}_{q}^{(\ell)}({\mathbf{W}_{j}}^{(\ell)}\odot\mathbf{m}_{q}^{(\ell)})+\mathbf{x}_{p}\widetilde{\mathbf{W}_{j}}^{(\ell)}\right)$ (11)
$\mathbf{h}_{u}^{(\ell+1)}=$ $\sigma\left(\sum_{v\in\mathcal{N}_{u}}{\tilde{a}_{uv}}\mathbf{h}_{v}^{(\ell)}({\mathbf{W}_{k}}^{(\ell)}\odot\mathbf{m}_{v}^{(\ell)})+\mathbf{x}_{u}\widetilde{\mathbf{W}_{k}}^{(\ell)}\right)$
where $\mathcal{N}_{p}$ and $\mathcal{N}_{u}$ respectively denote the neighbor sets of node $p$ and node $u$, $\tilde{a}_{pq}$ and $\tilde{a}_{uv}$ represent the adjacency coefficients between the node pairs $(p,q)$ and $(u,v)$, $\mathbf{m}_{q}$ and $\mathbf{m}_{v}$ are the modulation vectors of node $q$ and node $v$, $\mathbf{x}_{p}$ and $\mathbf{x}_{u}$ respectively represent the initial input features of node $p$ and node $u$. Thus, the equation of the $\ell+1$ layer feature is reformulated as
@@
4 Experiments
In this section, we first introduce the benchmark datasets, Human3.6M [30] and MPI-INF-3DHP [31], and the evaluation metrics used to assess the proposed framework, along with the experimental setup. We then evaluate the framework and compare it with current state-of-the-art methods, presenting both quantitative and qualitative results. Finally, we conduct ablation studies to assess the impact of each proposed design.
4.1 Experimental Setup
@@
Human3.6M [30] is a widely used dataset for 3D human pose estimation, comprising 3.6 million poses captured using RGB and ToF cameras from multiple viewpoints in real-world settings. Following the standard protocol, we use five subjects (S1, S5, S6, S7, S8) for training and two subjects (S9, S11) for evaluation. We evaluate performance using Mean Per Joint Position Error (MPJPE) under two protocols: Protocol 1 (P1), which aligns the root joints before calculating MPJPE, and Protocol 2 (P2), which involves Procrustes-MPJPE (PA-MPJPE), where the ground truth and the predicted pose are aligned through a rigid transformation.
MPI-INF-3DHP [31] is a large-scale dataset with over 2K videos featuring joint annotations of 13 keypoints in outdoor scenes, designed for both 2D and 3D human pose estimation. The ground truth is obtained through a multi-camera setup and a marker-less MoCap system. The evaluation metrics include MPJPE, Percentage of Correct Keypoints (PCK) at a threshold of 150 mm, and Area Under the Curve (AUC).
@@
We evaluate our model by comparing it with state-of-the-art video-based 3D human pose estimation methods, including transformer-based approaches such as StridedFormer [22], MHFormer [21], MixSTE [20], P-STMO [41], HSTFormer [18], Einfalt et al. [19], and PoseFormerV2 [16], as well as GCN-based methods like UGCN [42] and GLA-GCN [11]. We also compare our model with hybrid methods combining graphs and transformers, such as STCFormer [25], Motionbert [43], and MotionAGFormer [44].
@@
For the experiments, we use three types of 2D pose sequences as input: estimated 2D poses from the pre-trained CPN [45], Stacked Hourglass [46], and real 2D poses (ground truth). The model is trained for 40 epochs using the Adam optimizer with a batch size of 128. The learning rate starts at 0.001 and decays by 0.97 per epoch. Our architecture is configured with L = 6 transformer layers, F = 256 feature embedding dimensions, H = 8 self-attention heads, and N = 2 MHR-GCN layers. The model is implemented using PyTorch (version 2.1.0) and runs on an RTX 4090 GPU with driver version 525.105.17 and CUDA version 12.0.
4.2 Results and Analysis